OpenStudy (anonymous):

log10 (x+6/x-3) = 1 find x

OpenStudy (anonymous):

\[\log_{10} \]

OpenStudy (lalaly):

10=x+6/x-3 10x-30=x+6 9x=36 x=4

OpenStudy (anonymous):

its log with a little 10 not log with a big 10 behind it.

OpenStudy (lalaly):

yeh i know that, the base is 10 so its 10^1 =x+6/x-3 and same as i did